Cochrane, Inc., is considering a new three-year expansion project that requires an initial fixed asset investment of $2,610,000. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $2,320,000 in annual sales, with costs of $1,310,000. Required: If the tax rate is 40 percent, what is the OCF for this project? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars (e.g., 1,234,567).)

OCF=
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(2,610,000/3)=870000

OCF=(Sales-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(2,320,000-1,310,000)(1-0.4)+(0.4*870000)

=$954,000